OpenStudy (anonymous):

solve the system of equations y=2x-10 and x-y=7

OpenStudy (anonymous):

By substituion: put y=2x-10 in x-y=7 x-(2x-10)=7 x-2x+10=7 -x=7-10 -x=-3 x=3 Put x=3 in y=2x-10 y=2*3-10 y=-4

OpenStudy (anonymous):

Graphically you can draw the lines and find the point of intersection.

OpenStudy (anonymous):

There is one more method called Elimination :)

OpenStudy (anonymous):

y=2x-10 can be changed into 2x-y=10 now subtract equation 2 from 1 2x-y=10 x-y=7 ------------ x=3 put value of x in first equation you will get value of y.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

shahzadjalbani gave elimination method :)

OpenStudy (anonymous):

ah i like that way better i think